Pistol Corporation purchased 100 percent ownership of Scope Products on January 1, 20X6, for $64,000, at which time Scope Products reported retained earnings of $14,000 and capital stock outstanding of $27,000. The differential was attributable to patents with a life of four years. Income and dividends of Scope Products were
 

YearNet IncomeDividends
20X6$20,000
$8,000
20X7
28,000

10,000
20X8
36,000

10,000


Required:
1. Prepare the equity method entries that Pistol should record to account for this investment in 20X6, 20X7, and 20X8. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1

To find the differential value to record the amortization you take the initial investment - the reported retained earnings - the capital stock outstanding and divide the outcome by the lifespan of the patents. 

(64000-14000-27000) / 4 = 5750

image.png

To find the balance of the investment account take the inital investment + the total of the net income - the total of the dividends received - (the yearly diff x the years recorded)

64000 + (20000 + 28000 + 36000) - (8000 + 10000 + 10000) - (5750 * 3) = 102750

image.png

answered by: Patrick J Boyd
Answer #2

To find the differential value to record the amortization you take the initial investment - the reported retained earnings - the capital stock outstanding and divide the outcome by the lifespan of the patents. 

(64000-14000-27000) / 4 = 5750

image.png

To find the balance of the investment account take the inital investment + the total of the net income - the total of the dividends received - (the yearly diff x the years recorded)

64000 + (20000 + 28000 + 36000) - (8000 + 10000 + 10000) - (5750 * 3) = 102750

image.png


source: McGraw-Hill Connect Christensen, Advanced Financial Accounting, 12e
answered by: Patrick J Boyd